# Large Deviations Principle for Isoperimetry and Its Equivalence to Nonlinear Log-Sobolev Inequalities

### v3

@@ Abstract.
The isoperimetric problem is a classic topic in geometric measure theory, yet critical questions regarding the characterization of optimal solutions—even asymptotically optimal ones—remain largely unresolved. In this paper, we investigate the large deviations asymptotics (governing sequences of exponentially small sets) for the isoperimetric problem on the product Riemannian manifold $\mathcal{M}^{n}$ endowed with the product probability measure $\nu^{\otimes n}$, where $\mathcal{M}$ is a Riemannian manifold satisfying the curvature-dimension condition $\mathrm{CD}(0,\infty)$. When the probability measure $\nu$} admits a finite moment generating function for squared distance, we establish an exact characterization of the large deviations asymptotics of the isoperimetric profile, which reveals a precise equivalence between this asymptotic isoperimetric inequality and an established strengthening of log-Sobolev inequality—namely, the nonlinear log-Sobolev inequality. It is observed that conditional typical sets (i.e., $\mathbf{f}$-coordinate half-spaces for some vector $\mathbf{f}$ of functions), a fundamental concept from information theory, form an asymptotically optimal solution to the isoperimetric problem. This class of subsets further yields an upper bound on the isoperimetric profile in the central limit regime (concerning constant-volume sets), expressed as the Gaussian isoperimetric profile times the square root of the spectral gap. Furthermore, we apply our results to establish quantitative relations among the optimal constants in isoperimetric, concentration and transport inequalities, additionally covering isoperimetry for subexponentially and superexponentially small sets. Our results provide a rigorous justification from the perspective of nonlinear log-Sobolev inequalities for why isoperimetric minimizers behave fundamentally differently across spaces with distinct geometric structures. We adopt a new proof idea in this paper which integrates tools from information theory, optimal transport, and geometric measure theory.
@@ 1. Introduction
Let $(\mathcal{M},g)$ be a finite-dimensional smooth complete oriented connected Riemannian manifold with induced geodesic distance $d$. Let $\nu=e^{-V}\mathrm{vol}$ with $V\in C^{2}(\mathcal{M})$ be a reference Borel probability measure on $\mathcal{M}$ that is absolutely continuous with respect to the Riemannian volume $\mathrm{vol}$. We consider the $n$-fold product space $\mathcal{M}^{n}$, equipped with the product measure $\nu_{n}:=\nu^{\otimes n}$ and the product distance $d_{n}(\mathbf{x},\mathbf{y})=\sqrt{\sum_{i=1}^{n}d(x_{i},y_{i})^{2}}$, where $\mathbf{x}=(x_{1},x_{2},...,x_{n})$.
@@
where $A$ is a Borel set. Note that there are several other ways in the literature to define the perimeter of a set, but they coincide when the set admits smooth boundary, and thus, lead to the same isoperimetric profile in our setup; see details in [10, 5, pp. 32-33].
@@ 1.1.1. Large Deviations Principle
Assumption 1 (Convexity).
A natural convexity condition on a manifold with density is the following curvature-dimension condition $\mathrm{CD}(0,\infty)$:
$\mathrm{Ric}_{g}+\mathrm{Hess}_{g}V\geq 0\;\;\text{ as 2-tensor fields},$
@@
Since $I_{n}(a)/a$ is non-increasing in $a$ under Assumption 1 [45, Proposition 3.1], $\Lambda$ is non-decreasing.
@@
Under Assumption 1, $\Lambda$ is non-decreasing.
Let $P_{2}^{\mathrm{ac}}(\mathcal{M})$ be the set of probability measures which are absolutely continuous with respect to $\nu$ and have finite second-order moments. We also need the concept of nonlinear log-Sobolev inequality and a generalized version of half-spaces.
@@
Given a probability measure $\nu$, for a convex function $\theta:[0,\infty]\to[0,\infty]$, the $\theta$-nonlinear log-Sobolev inequality is defined to be the following inequality
(1.6) $\theta(D(\mu\|\nu))\leq I(\mu\|\nu),\;\forall\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M}),$
@@
$D(\mu\|\nu):=\int\rho\ln\rho d\nu\;\textrm{ and }\;I(\mu\|\nu):=\int\frac{|\nabla\rho|^{2}}{\rho}d\nu$
@@
The inequality in (1.6) holds for all $\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M})$ if and only if it holds for all $\mu\in P^{\mathrm{ac}}(\mathcal{M})$ (the set of probability measures which are absolutely continuous with respect to $\nu$) [54].
The nonlinear log-Sobolev inequality, also known as the entropy-energy inequality or the entropy isoperimetric inequality, can be seen as a strengthening of the standard (linear) log-Sobolev inequality since it provides more information about the tradeoff between entropy and Fisher information. This concept was investigated in several important works in the literature, e.g., Costa and Cover [20], F. Y. Wang [66], Bakry, Gentil, and Ledoux [7], as well as Polyanskiy and Samorodnitsky [55].
Any convex $\theta$ that admits a $\theta$-nonlinear log-Sobolev inequality must satisfy that $\theta(0)=0$ and $\theta$ is non-decreasing. In particular, when $\theta(t)=Kt$ for some constant $K>0$, then the $\theta$-nonlinear log-Sobolev inequality reduces to the standard log-Sobolev inequality. Define
$\Theta(\alpha):=\inf_{\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M}):D(\mu\|\nu)\geq\alpha}I(\mu\|\nu).$
Then, given a probability measure $\nu$, the optimal convex function $\theta$ in the $\theta$-nonlinear log-Sobolev inequality in (1.6) is the lower convex envelope of $\Theta$, given by
@@
(1.7) $=\inf_{\begin{subarray}{c}\lambda\in[0,1],\mu_{1},\mu_{2}\in P_{2}^{\mathrm{ac}}(\mathcal{M}):\\ \lambda D(\mu_{1}\|\nu)+(1-\lambda)D(\mu_{2}\|\nu)\geq\alpha\end{subarray}}\lambda I(\mu_{1}\|\nu)+(1-\lambda)I(\mu_{2}\|\nu).$
@@
The second concept we need is a generalized version of half-spaces which incorporates many important concepts as special cases, e.g., the concept of typical sets from information theory.
@@
The concept of $\mathbf{f}$-coordinate half-space reduces to the Orlicz ball if $f_{i}$ are Young functions (and t>0-t>0). However, we do not pose any restriction on functions $f_{i}$ here (except for measurability), and thus our concept is much more general and incorporates standard half-spaces as special cases.
@@
Here, $(\mu_{i})_{i\in[n]}$ can be seen as a conditional probability measure $\mu_{X|W}$ with $\mu_{X|W=i}=\mu_{i}$, $\frac{1}{n}\sum_{i=1}^{n}\ln\frac{{\text{\rm d}}\mu_{i}}{{\text{\rm d}}\nu}(x_{i})$ is known as the conditional information density of $\mu_{X|W}$ w.r.t. $\nu$ given $\mu_{W}\sim\mathrm{Unif}[n]$, and its expectation is $D(\mu_{X|W}\|\nu|\mu_{W}):=\frac{1}{n}\sum_{i=1}^{n}D(\mu_{i}\|\nu)$ which is known as the conditional relative entropy. That is why it is called the conditional typical set. In addition, the standard definition of the conditional relative entropy typical set is $\big\{\mathbf{x}\in\mathcal{M}^{n}:\big|\sum_{i=1}^{n}\ln\frac{{\text{\rm d}}\mu_{i}}{{\text{\rm d}}\nu}(x_{i})-\sum_{i=1}^{n}D(\mu_{i}\|\nu)\big|\leq n\epsilon\big\}$ [21, (11.206)]. Our definition here is slightly different from the standard one, and has benefit that it incorporates the isoperimetric minimizers for Gaussian and Lebesgue measures as special cases, which is explained below.
@@
The concept of $\mathbf{f}$-coordinate half-space generalizes the standard half-spaces $\{\mathbf{x}\in\mathbb{R}^{n}:\sum_{i=1}^{n}w_{i}x_{i}\leq t\}$ in Euclidean spaces, and the former reduces to the latter when $f_{i}(x_{i})=-w_{i}x_{i}$ and $t$ is rechosen, or essentially equivalently, set $\nu=\mathcal{N}(0,1),\mu_{i}=\mathcal{N}(w_{i},1)$ in the conditional typical set. It also generalizes the Euclidean balls $\left\{\mathbf{x}\in\mathbb{R}^{n}:\sum_{i=1}^{n}x_{i}^{2}\leq t\right\}$ with $t>0$, and the former reduces to the latter when $f_{i}(x_{i})=-x_{i}^{2}$ and $t$ is rechosen, or essentially equivalently, set $\nu=\mathrm{Lebesgue},\mu=\mathcal{N}(0,c)$ with $c>0$ in the conditional typical set (although the Lebesgue measure is an infinite measure and not considered in this paper). Note that standard half-spaces and Euclidean balls are respectively isoperimetric minimizers for Gaussian measures and Lebesgue measures.
@@
We now establish the large deviations principle for the isoperimetric problem, a result of particular interest as it reveals a precise equivalence between the asymptotic isoperimetric inequality and the nonlinear log-Sobolev inequality. We also need the following assumption.
Assumption 2 (Full Exponential Integrability of Squared Distance).
For some $x_{0}\in\mathcal{M}$ (equivalently for any $x_{0}\in\mathcal{M}$),
$\mathbb{E}_{\nu}[e^{\lambda d^{2}(X,x_{0})}]<\infty,\;\forall\lambda\in(0,\infty).$
In other words, the probability measure $\nu$} admits a finite moment generating function for squared distance.
This assumption is satisfied if $\mathcal{M}$ is a bounded Riemann manifold or if $\mathcal{M}$ is a Euclidean space with $\nu$ having density $\propto e^{-C|x|^{\beta}}$ where $C>0$ and $\beta>2$.
@@
Under Assumptions 1 and 2, it holds that for any $\alpha>0$,
@@
Moreover, a sequence of conditional typical sets $A_{n,\epsilon}((\mu_{1}^{*},...,\mu_{1}^{*},\mu_{2}^{*},...,\mu_{2}^{*})\|\nu)$ with $\mu_{1}^{*},\mu_{2}^{*}$ respectively appearing $\lfloor n\lambda^{*}\rfloor$ and $n-\lfloor n\lambda^{*}\rfloor$ times asymptotically attains the limit $\Lambda(\alpha)$ as $n\to\infty$ first and $\epsilon\to 0$ then, where $(\lambda^{*},\mu_{1}^{*},\mu_{2}^{*})$ is the optimal solution in (1.7). In addition, the inequality $\overline{\Lambda}(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}$} holds even without Assumption 2.
@@
Although this theorem does not include the Gaussian measure case and the Lebesgue measure case, for these two cases, the conditional typical sets with $\mu$ set to the extremizers (see [55]) respectively reduce to standard half-spaces and Euclidean balls, consistent with the already known fact that these two kinds of sets are isoperimetric extremizers for these two cases. Moreover, by this already known fact, Equation (1.9) still holds for these two cases, suggesting that Theorem 1 might hold for a much larger class of measures.
This theorem can be alternatively formulated as follows: For any convex $\theta:[0,\infty)\to[0,\infty)$,
${\displaystyle\Lambda(\alpha)\geq\sqrt{\theta(\alpha)}},\;\forall\alpha>0\quad\Longleftrightarrow\quad I(\mu\|\nu)\geq\theta(D(\mu\|\nu)),\;\forall\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M}).$
Equation (1.9) is also equivalent to that for any $\alpha>0$,
@@
where $a_{n}=e^{-n\alpha}$ and $I_{\mathrm{G}}$ is the isoperimetric profile for the standard Gaussian measure. That is, as the volume vanishes exponentially, the asymptotic behavior of the isoperimetric profile for any product probability measure resembles the one for the standard Gaussian measure, but with different factors that are determined by the exponential convergence rate of the volume. One prototypical case occurs when $\Theta(\alpha)$ grows linearly in $\alpha$—for instance, $\Theta(\alpha)=2\alpha$ for Gaussian measures—implying that $I_{n}(a)$ behaves as $C\cdot a\sqrt{\ln\frac{1}{a}}$, with standard half-spaces likely serving as approximate extremizers for small volumes. A second prototypical case is when $\Theta(\alpha)$ grows exponentially in $\alpha$, as illustrated in the subsequent example, which implies that $I_{n}(a)$ behaves as $C\cdot\sqrt{n}\,a^{1-1/n}$, with geodesic balls acting as approximate extremizers for small volumes. Theorem 1 provides a rigorous justification from the perspective of nonlinear log-Sobolev inequalities for why isoperimetric minimizers of small volumes behave fundamentally differently across spaces with distinct geometric structures.
Example 1 (Exponential Growth of $\Theta$).

### v4

@@ Abstract.
The isoperimetric problem is a classic topic in geometric measure theory, yet critical questions regarding the characterization of optimal solutions—even asymptotically optimal ones—remain largely unresolved. In this paper, we investigate the large deviations asymptotics (governing sequences of exponentially small sets) for the isoperimetric problem on the product Riemannian manifold $\mathcal{M}^{n}$ endowed with the product probability measure $\nu^{\otimes n}$, where $(\mathcal{M},\nu)$ is a weighted Riemannian manifold with nonnegative Bakry–Émery–Ricci curvature. We establish an exact characterization of the large deviations asymptotics of the isoperimetric profile, which reveals a precise equivalence between this asymptotic isoperimetric inequality and an established strengthening of log-Sobolev inequality—namely, the nonlinear log-Sobolev inequality. It is observed that conditional typical sets (i.e., $\mathbf{f}$-coordinate half-spaces for some vector $\mathbf{f}$ of functions), a fundamental concept from information theory, form an asymptotically optimal solution to the isoperimetric problem. This class of subsets further yields an upper bound on the isoperimetric profile in the central limit regime (concerning sets of constant weighted volume). Although our results are stated for product spaces, they imply certain isoperimetric inequalities for non-product spaces, e.g., they can be used to recover the weaker equivalence established by Ledoux and Bobkov for arbitrary non-product spaces or used to establish quantitative relations among the optimal constants in isoperimetric, concentration and transport inequalities for product or non-product spaces. Our results provide a rigorous justification from the perspective of nonlinear log-Sobolev inequalities for why isoperimetric minimizers behave fundamentally differently across spaces with distinct geometric structures. Our proof idea is a new framework which integrates tools from information theory, optimal transport, and geometric measure theory.
@@ 1. Introduction
Let $(\mathcal{M},g)$ be a $k$-dimensional smooth complete oriented connected Riemannian manifold with induced geodesic distance $d$. We assume $k\geq 1$ and $(\mathcal{M},g)=(\mathbb{R},|\cdot|)$ if $k=1$. Let $\nu=e^{-V}\mathrm{vol}$ with $V\in C^{2}(\mathcal{M})$ be a reference Borel probability measure on $\mathcal{M}$ that is absolutely continuous with respect to the Riemannian volume $\mathrm{vol}$. We consider the $n$-fold product space $\mathcal{M}^{n}$, equipped with the product measure $\nu_{n}:=\nu^{\otimes n}$ and the product distance $d_{n}(\mathbf{x},\mathbf{y})=\sqrt{\sum_{i=1}^{n}d(x_{i},y_{i})^{2}}$, where $\mathbf{x}=(x_{1},x_{2},...,x_{n})$.
@@
where $A$ is a Borel set. Note that there are several other ways in the literature to define the perimeter of a set, but they coincide when the set admits smooth boundary, and thus, lead to the same isoperimetric profile in our setup; see details in [15, 5, pp. 32-33].
@@ 1.1.1. Large Deviations Principle
A natural convexity condition on a manifold with density is the curvature-dimension condition $\mathrm{CD}(0,\infty)$.
Condition 1 (Nonnegative Bakry–Émery–Ricci Curvature).
We assume that
$\mathrm{Ric}_{V}:=\mathrm{Ric}_{g}+\mathrm{Hess}_{g}V\geq 0\;\;\text{ as 2-tensor fields},$
@@
Since $I_{n}(a)/a$ is non-increasing in $a$ under Condition 1 (see e.g. [52, Proposition 3.1]), $\Lambda$ is non-decreasing.
@@
Under Condition 1, $\Lambda$ is non-decreasing.
Let $\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M})$ be the set of probability measures which are absolutely continuous with respect to $\nu$ and have finite second-order moments. We also need the concept of nonlinear log-Sobolev inequality and a generalized version of half-spaces.
@@
For a convex function $\theta:[0,\infty]\to[0,\infty]$, we say a probability measure $\nu$ admits the $\theta$-nonlinear log-Sobolev inequality if
(1.6) $\theta(D(\mu\|\nu))\leq I(\mu\|\nu),\;\forall\mu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}),$
@@
$D(\mu\|\nu):=\int\rho\ln\rho{\text{\rm d}}\nu\;\textrm{ and }\;I(\mu\|\nu):=\int\frac{|\nabla\rho|^{2}}{\rho}{\text{\rm d}}\nu$
@@
The inequality in (1.6) holds for all $\mu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M})$ if and only if it holds for all $\mu\in\mathscr{P}^{\mathrm{ac}}(\mathcal{M})$ (the set of probability measures which are absolutely continuous with respect to $\nu$) [58].
The nonlinear log-Sobolev inequality, also known as the entropy-energy inequality or the entropy isoperimetric inequality, can be seen as a strengthening of the standard (linear) log-Sobolev inequality since it provides more information about the tradeoff between entropy and Fisher information. This concept was investigated in several important works in the literature, e.g., Costa and Cover [26], Bakry [9], F. Y. Wang [75], Bakry, Gentil, and Ledoux [10], as well as Polyanskiy and Samorodnitsky [59].
For a convex $\theta$, a $\theta$-nonlinear log-Sobolev inequality holds only if $\theta(0)=0$ and $\theta$ is non-decreasing. In particular, when $\theta(t)=Kt$ for some constant $K>0$, then the $\theta$-nonlinear log-Sobolev inequality reduces to the standard log-Sobolev inequality. Define
$\Theta(\alpha):=\inf_{\mu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}):D(\mu\|\nu)\geq\alpha}I(\mu\|\nu).$
Then, given the probability measure $\nu$, the optimal convex function $\theta$ in the $\theta$-nonlinear log-Sobolev inequality in (1.6) is the lower convex envelope of $\Theta$, given by
@@
(1.7) $=\inf_{\begin{subarray}{c}\lambda\in[0,1],\mu_{1},\mu_{2}\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}):\\ \lambda D(\mu_{1}\|\nu)+(1-\lambda)D(\mu_{2}\|\nu)\geq\alpha\end{subarray}}\lambda I(\mu_{1}\|\nu)+(1-\lambda)I(\mu_{2}\|\nu).$
@@
The second concept we need is a generalized version of half-spaces which incorporates several important concepts as special cases, e.g., the concept of typical sets from information theory.
@@
The concept of $\mathbf{f}$-coordinate half-space reduces to the Orlicz ball if $f_{i}$ are Young functions (and t>0-t>0). However, we do not pose any restriction on functions $f_{i}$ here (except for measurability), and thus our concept is more general and incorporates standard half-spaces as special cases.
@@
Here, $(\mu_{i})_{i\in[n]}$ can be seen as a conditional probability measure $\mu_{X|W}$ with $\mu_{X|W=i}=\mu_{i}$, $\frac{1}{n}\sum_{i=1}^{n}\ln\frac{{\text{\rm d}}\mu_{i}}{{\text{\rm d}}\nu}(x_{i})$ is known as the conditional information density of $\mu_{X|W}$ w.r.t. $\nu$ given $\mu_{W}\sim\mathrm{Unif}[n]$, and its expectation $D(\mu_{X|W}\|\nu|\mu_{W}):=\frac{1}{n}\sum_{i=1}^{n}D(\mu_{i}\|\nu)$ is known as the conditional relative entropy. That is why it is called the conditional typical set. In addition, the standard definition of the conditional relative entropy typical set is $\big\{\mathbf{x}\in\mathcal{M}^{n}:\big|\sum_{i=1}^{n}\ln\frac{{\text{\rm d}}\mu_{i}}{{\text{\rm d}}\nu}(x_{i})-\sum_{i=1}^{n}D(\mu_{i}\|\nu)\big|\leq n\epsilon\big\}$ [27, (11.206)]. Our definition here is slightly different from the standard one, and has benefit that it incorporates the isoperimetric minimizers for Gaussian and Lebesgue measures as special cases, which is explained below.
@@
The concept of $\mathbf{f}$-coordinate half-space generalizes the standard half-spaces $\{\mathbf{x}\in\mathbb{R}^{n}:\sum_{i=1}^{n}w_{i}x_{i}\leq t\}$ in Euclidean spaces, and the former reduces to the latter when $f_{i}(x_{i})=-w_{i}x_{i}$ and $t$ is rechosen properly, or essentially equivalently, set $\nu=\mathcal{N}(0,1),\mu_{i}=\mathcal{N}(w_{i},1)$ in the conditional typical set. It also generalizes the Euclidean balls $\left\{\mathbf{x}\in\mathbb{R}^{n}:\sum_{i=1}^{n}x_{i}^{2}\leq t\right\}$ with $t>0$, and the former reduces to the latter when $f_{i}(x_{i})=-x_{i}^{2}$ and $t$ is rechosen properly, or essentially equivalently, set $\nu=\mathrm{Lebesgue},\mu=\mathcal{N}(0,c)$ with $c>0$ in the conditional typical set (although as an infinite measure, the Lebesgue measure is not considered in this paper). In both these two cases, $\mu_{i}$ are the extremizers in the optimal nonlinear log-Sobolev inequality for $\nu$ [9, 59]. Furthermore, it is well known that standard half-spaces and Euclidean balls are respectively isoperimetric minimizers for Gaussian measures and Lebesgue measures. This observation indicates that the isoperimetric minimizers perhaps can be identified by the extremizers in the optimal nonlinear log-Sobolev inequality.
@@
We now formalize the above observation and establish the large deviations principle for the isoperimetric problem, a result of particular interest as it reveals a precise equivalence between the asymptotic isoperimetric inequality and the nonlinear log-Sobolev inequality. We state our results by splitting into cases according to whether the following condition holds.
Condition 2 (Exponential Integrability).
For some (thus all) $x_{0}\in\mathcal{M}$,
$\mathbb{E}_{\nu}[e^{\lambda d^{2}(X,x_{0})}]<\infty,\;\exists\lambda\in(0,\infty).$
This condition can be partitioned into two cases:
(1.9) $\mathbb{E}_{\nu}[e^{\lambda d^{2}(X,x_{0})}]<\infty,\;\forall\lambda\in(0,\infty),$
and
(1.10) $\mathbb{E}_{\nu}[e^{\lambda_{1}d^{2}(X,x_{0})}]<\mathbb{E}_{\nu}[e^{\lambda_{2}d^{2}(X,x_{0})}]=\infty,\;\exists 0<\lambda_{1}<\lambda_{2}<\infty.$
Under Condition 1, Condition 2 is in fact equivalent to that $\nu$ admits a standard log-Sobolev inequality [74]. We need the following assumption when the condition in (1.10) holds.
Condition 3 (Ricci Curvature Bounded Below).
$\mathrm{Ric}\geq K$ for some $K\in\mathbb{R}$.
@@
Suppose the weighted Riemannian manifold $(\mathcal{M},d,\nu)$ satisfies Condition 1.
  1. (1)
    Suppose Condition 2 holds. Suppose Condition 3 (or Conditions 4, 5, or 6 given in Section 3.3) holds when the condition in (1.10) holds. Then, for any $\alpha\geq 0$,
@@
    Moreover, a sequence of conditional typical sets $A_{n,\epsilon}((\mu_{1}^{*},...,\mu_{1}^{*},\mu_{2}^{*},...,\mu_{2}^{*})\|\nu)$ with $\mu_{1}^{*},\mu_{2}^{*}$ respectively appearing $\lfloor n\lambda^{*}\rfloor$ and $n-\lfloor n\lambda^{*}\rfloor$ times asymptotically attains the limit $\Lambda(\alpha)$ as $n\to\infty$ first and $\epsilon\to 0$ then, where $(\lambda^{*},\mu_{1}^{*},\mu_{2}^{*})$ is the optimal solution in (1.7).
  2. (2)
    If Condition 2 does not hold, then, for any $\alpha\geq 0$,
    $\Lambda(\alpha)=\sqrt{\breve{\Theta}(\alpha)}=0$}.
Although the theorem above is valid for spaces that are general enough, it does not cover isoperimetric problems for log-concave probability measures defined on closed convex subsets of Euclidean spaces. To cover them, we extend the theorem above to an $\mathrm{RCD}(0,\infty)$-subspace of the Riemannian manifold $\mathcal{M}$. Here, an $\text{RCD}(0,\infty)$-space (a space satisfying Riemannian Ricci Curvature Dimension bound $\text{RCD}(0,\infty)$) is a generalization of Riemannian manifolds satisfying Condition 1 to possibly non-smooth setting, e.g., compact subsets of Euclidean spaces equipped with log-concave probability measures. From the perspective of the geometry of optimal transport, an $\text{RCD}(0,\infty)$-space is a Polish metric probability measure space such that infinitesimal Hilbertianity is satisfied (i.e., the Cheeger energy is a quadratic form), and meanwhile, the relative entropy functional $D(\cdot\|\nu)$ is not merely $K$-geodesically convex on the Wasserstein space $(\mathscr{P}_{2}(\mathcal{X}),\mathsf{W}_{2})$, but that this convexity is robust: it extends to hold along all interpolated measures induced by weighted optimal geodesic plans (or test plans). See details for the definition of the $\text{RCD}(0,\infty)$-space, or equivalently, the Riemannian Energy measure space satisfying $\text{BE}(K,\infty)$ in Section 3.2 or [7, 8].
Theorem 2.
Let $\nu$ be an absolutely continuous probability measure with density $f$ (w.r.t. Riemannian volume) on a finite-dimensional smooth complete oriented connected Riemannian manifold $\mathcal{M}$ such that $\nu$ together with its support forms a $\mathrm{RCD}(0,\infty)$-subspace of $\mathcal{M}$. Suppose that there exists a sequence of probability measures $\nu_{m}\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}),m\in\mathbb{N}$ with densities $f_{m}$ (w.r.t. Riemannian volume) satisfying Condition 1 and $f_{m}\geq(1-1/m)f$, as well as Condition 3 if the condition in (1.10) holds. Then, for any $\alpha\geq 0$, $\Lambda(\alpha)=\sqrt{\breve{\Theta}(\alpha)}$}.
@@
The convergence of $\{\nu_{m}\}$ here is called “convergence from above” by E. Milman [51]. In addition, another notion, “convergence from within”, was also introduced by him. Theorem 2 still holds if convergence of $\{\nu_{m}\}$ from above is replaced by convergence from within.
The equivalence $\Lambda=\sqrt{\breve{\Theta}}$ in the two theorems above can be alternatively formulated as follows: For any convex $\theta:[0,\infty)\to[0,\infty)$,
${\displaystyle\Lambda(\alpha)\geq\sqrt{\theta(\alpha)}},\;\forall\alpha\geq 0\quad\Longleftrightarrow\quad I(\mu\|\nu)\geq\theta(D(\mu\|\nu)),\;\forall\mu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}).$
It is also equivalent to that for any $\alpha>0$,
@@
where $a_{n}=e^{-n\alpha}$ and $I_{\mathrm{G}}$ is the isoperimetric profile for the standard Gaussian measure. That is, as the weighted volume vanishes exponentially, the asymptotic behavior of the isoperimetric profile for any product probability measure resembles the one for the standard Gaussian measure, but with different factors that are determined by the exponential convergence rate of the weighted volume. One prototypical case occurs when $\Theta(\alpha)$ grows linearly in $\alpha$—for instance, $\Theta(\alpha)=2\alpha$ for Gaussian measures—implying that $I_{n}(a)$ behaves as $C\cdot a\sqrt{\ln\frac{1}{a}}$, with standard half-spaces likely serving as approximate extremizers for small weighted volumes. A second prototypical case is when $\Theta(\alpha)$ grows exponentially in $\alpha$, as illustrated in the subsequent example, which implies that $I_{n}(a)$ behaves as $C\cdot\sqrt{n}\,a^{1-1/n}$, with geodesic balls acting as approximate extremizers for small weighted volumes. Theorem 1 provides a rigorous justification from the perspective of nonlinear log-Sobolev inequalities for why isoperimetric minimizers of small weighted volumes behave fundamentally differently across spaces with distinct geometric structures.
Example 1 (Log-Concave Probability Measures).
All smooth log-concave probability densities on Euclidean spaces satisfy Conditions 1 and 3. Moreover, by standard regularization techniques (e.g., the Moreau-Yosida regularization and convolution with Gaussian kernels), all log-concave probability densities defined on closed convex subsets of Euclidean spaces can be approximated by smooth log-concave probability densities defined on the whole Euclidean spaces in the sense as in Theorem 2. Moreover, closed convex subsets equipped with log-concave probability densities satisfy $\mathrm{RCD}(0,\infty)$. Thus, $\Lambda=\sqrt{\breve{\Theta}}$ holds for all (smooth or non-smooth) log-concave probability measures on Euclidean spaces (supported on the whole spaces or convex subsets). The standard Gaussian measure is one of the most important examples, which admits $\breve{\Theta}(\alpha)=\Theta(\alpha)=2\alpha$ and thus, $\Lambda(\alpha)=\sqrt{2\alpha}$, consistent with the formula in (1.4).
Example 2 (Bounded Manifolds admit Exponential Growth of $\Theta$).